OpenStudy (anonymous):

Hi, I need help on a factoring problem. x^3=27

OpenStudy (curry):

x= 3 dont forgot medal

OpenStudy (anonymous):

Don't forget what?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

How to I factor it?

OpenStudy (curry):

wat to the power of 3 equals 27

OpenStudy (anonymous):

Oops! I typed it wrong, I need to factor x^3+27

OpenStudy (curry):

so in other words there are 2 equations to this

OpenStudy (curry):

oh ok that is more like it

OpenStudy (anonymous):

One way to do this is to take log base 3 on both sides. You get 3logx=log27 =3logx=3 logx=1 Since its base 3 x=3^1=3

OpenStudy (anonymous):

oh in that case, x=-3

OpenStudy (anonymous):

What's log base?

OpenStudy (anonymous):

I need to factor x^3+27

OpenStudy (anonymous):

Okay,lets try it differently, x^3-3x^2+3x^2-9x+9x+27= (x+3)(x^2-3x+9)=0 This gives x=-3 as the only real root.

OpenStudy (anonymous):

Ok, thank you!